A man can throw a ball up to 80m the maximum height upto which he can rise the ball
 

Grade:11

1 Answers

Khimraj
5 years ago
R = u2 * Sin 2θ / g.
As, the range is maximum, so, θ = 450
So, 80 = u2 /g.
Now, H =  u2 Sin2θ / 2g.
Putting value of θ = 45, u2 / g = 80, we get
H = 80/4  = 20m
Hope it clears. If you like answer then please approve it.